OpenStudy (anonymous):

cos(4x)cos(2x)-sin(2x)sin(4x) is equivalent to which of the following? cos(6x) or sin(6x) or sin(2x) or none of these or cos(2x) i think it's cos(6x) but not sure

OpenStudy (anonymous):

Recall the double-angle formulas.\[\sin(A\pm B)=\sin A \cos B \pm \cos B \sin A\]\[\cos(A\pm B)=\cos A \cos B \mp \sin A \sin B\]PS - you are correct. :P